# Exact treatment of the quantum Langevin equation under time-dependent system-bath coupling via a train of delta distributions

### v1

@@ A.1 Transformation of covariance matrices
As described in Sec. III.3, Gaussian states are fully determined by the first moment $\bar{\bm{R}}\equiv\braket{\bm{R}}$ and the covariance matrix $\mathbb{V}$. Any tensor product of Gaussians is Gaussian, and Gaussian operations are those CP maps preserving the Gaussianity. In particular, the full Hamiltonian in the Caldeira-Leggett model generates a Gaussian unitary. If the initial joint state is a product of Gaussian states, $\rho_{\text{tot}}(0)=\rho_{\text{s}}(0)\otimes\rho_{\text{E}}(0)$, then the final reduced state of the system, $\rho_{\text{s}}(t)=\operatorname{Tr}_{\text{E}}[U_{t}\rho_{\text{tot}}(0)U_{t}^{\dagger}]$, after the Gaussian unitary time-evolution $U_{t}\equiv U(t,0)$ remains Gaussian.
@@
A.2 Transformation of covariance matrices under the Markovian dynamics
Although Eq. (57) does not hold in general, we will show that it does for Markovian dynamics. To this end, we introduce the concepts of universal dynamical maps (UDMs) and Markovian dynamical maps.
The quantum Markovianity of a quantum dynamical map is often characterized by the completely positive trace-preserving (CPTP) property, along with the phenomenologically motivated semigroup property, which ensures an irreversible dynamical process. These properties are equivalent to the dynamical map satisfying the well-known GKSL master equation [2], [^6] The CPTP and semigroup properties are equivalent to the generalized GKSL master equation, where the Hamiltonian, coupling, and Kraus operators can depend on time. forming the foundation of Markovian quantum dynamics.
In this work, we define Markovian dynamics as those that can be expressed as a CPTP semigroup map when the initial condition is a Gaussian state. However, we emphasize that this is not the only definition of Markovianity, as various alternative characterizations exist [25].
Let us begin by defining Markovianity [26]. Consider a quantum channel (i.e., a CPTP map) that maps a system’s arbitrary quantum state from $t=0$ to time $t$, $\mathcal{E}_{t,0}:\rho_{\text{s}}(0)\mapsto\rho_{\text{s}}(t)$. Quantum channels can be utilized even when a system is initially entangled with an environment. A universal dynamical map (UDM) is a special class of quantum channels, where the system’s input state is assumed to be uncorrelated with the environment. For example, a map induced by $\rho_{\text{s}}(t)=\operatorname{Tr}_{\text{E}}[U_{t}\rho_{\text{tot}}(0)U_{t}^{\dagger}]$ is a UDM if the system is initially uncorrelated with the environment.
A Markovian dynamics can be characterized in terms of the divisibility of a UDM. Suppose a map $\mathcal{E}_{t,0}$ is a UDM and it can be decomposed into two maps:
In general, however, it is not necessarily true that both $\mathcal{E}_{t,s}$ and $\mathcal{E}_{s,0}$ are UDMs. For example, $\mathcal{E}_{s,0}$ can be a UDM, but $\mathcal{E}_{t,s}$ is generally not, as the first UDM $\mathcal{E}_{s,0}$ entangles the system and the environment. A UDM $\mathcal{E}_{t,0}$ is called a Markovian dynamical map if both maps $\mathcal{E}_{t,s}$ and $\mathcal{E}_{s,0}$ are UDMs. This is also known as the one-parameter semigroup property, and it can be shown that it is equivalent to the generalized GKSL equation [26].
Next, we consider how the above Markovianity condition applies to the formulation using covariance matrices in Gaussian states. In particular, we show that if the time-evolution is described by a Markovian dynamical map, then the matrices $\mathbb{T}$ and $\mathbb{N}$ satisfy
@@ A.2 Transformation of covariance matrices under the Markovian dynamics
where $\mathbb{T}_{t-s}\equiv\mathbb{T}(t,s)$ and $\mathbb{N}_{t-s}\equiv\mathbb{N}(t,s)$ are the matrices associated with a UDM $\mathcal{E}_{t,s}$.
@@
where $\mathcal{E}_{t,0}$ is a UDM and we used the relation (33). Suppose the UDM $\mathcal{E}_{t,0}$ is a Markovian dynamical map. Then, we have the following relations:
@@
Appendix B Proof of Markovianity in the Limit $\Lambda T/N\gg 1$ and $\gamma/\Omega\ll 1$
In this section, we demonstrate that the two approximations, $\Lambda T/N\gg 1$ and $\gamma/\Omega\ll 1$, imply the Markovianity condition (59). We refer to these as the Markov and Born approximations, respectively.
@@ Appendix B Proof of Markovianity in the Limit $\Lambda T/N\gg 1$ and $\gamma/\Omega\ll 1$
where the functions $G_{f_{Q}}(t)$ and $G_{f_{P}}(t)$ are defined in Eq. (24), and the matrix components $N_{QQ}(t)$, $N_{QP}(t)$, $N_{PQ}(t)$, and $N_{PP}(t)$ are specified in Eq. (38).
Next, we recall that the quantity $G_{f}(t)$ can be expressed as a combination of the function $f$ and a matrix polynomial in $K$, as defined in Eq. (18). Since the matrix $K$ is of order $\mathcal{O}(\gamma/\Omega)$, the first-order Born approximation of $G_{f}(t)$ is given by
$G_{f}(t)=f(t)+\sum_{l=1}^{N-1}K_{t,t_{l}}f(t_{l})+\mathcal{O}\left((\gamma/\Omega)^{2}\right),$ (63)
for $f=f_{Q}$ and $f=f_{P}$. Note that the term with $l=N$ does not contribute to the sum above, since $K_{t,t_{N}}=0$ when $t=t_{N}$.
The explicit form of $K_{t,t_{l}}$ is given by
$K_{t,t_{l}}=$ $\sum_{k=1}^{N}\gamma\left(\frac{\Lambda T}{N}\right)^{2}e^{-\frac{\Lambda T}{N}|k-l|}\Theta(t_{k}-t_{l})$
×$\times\frac{\sin\left[\Omega(t-t_{k})\right]}{\Omega}\Theta(t-t_{k}).$ (64)
Under the Markov approximation, $\Lambda T/N\gg 1$, the leading-order term in the above summation is $k=l+1$:
$K_{t,t_{l}}=$ $\gamma\left(\frac{\Lambda T}{N}\right)^{2}e^{-\Lambda\frac{T}{N}}\frac{\sin\left[\Omega(t-t_{l+1})\right]}{\Omega}$
$+\mathcal{O}\left(\left(\frac{\Lambda T}{N}\right)^{2}\left(e^{-\frac{\Lambda T}{N}}\right)^{2}\right).$ (65)
Thus, under the combined Born-Markov approximation, the function $G_{f}(t)$ becomes
$G_{f}(t)\simeq$ $f(t)+\sum_{l=1}^{N-1}\frac{T^{2}}{N^{2}}\Gamma\left(\frac{T}{N}\right)\frac{\sin\left[\Omega(t-t_{l+1})\right]}{\Omega}f(t_{l})\,,$ (66)
where $\Gamma(t)$ denotes the Lorentz-Drude spectral density defined in Eq. (46).
This expression indicates that, under the Born-Markov approximation, the leading-order contribution arises from propagation over a single time step interval, i.e., from $t_{l}$ to $t_{l+1}$. Propagations over longer intervals, such as from $t_{l}$ to $t_{l+n}$ for $n=2,3,\dots$, are exponentially suppressed by terms of the form $(\Lambda T/N)^{2}\exp(-n\Lambda T/N)$, where $n$ is the number of time steps. Therefore, retaining only the one-time-step propagation diagram, as illustrated in Fig. 2, corresponds to the Markovian dynamics.
We now proceed to verify the semigroup property $\mathbb{T}_{t}=\mathbb{T}_{t-s}\mathbb{T}_{s}$. Since $t>s>0$, without loss of generality, let us denote $s=t_{j}$ for some $N>j>0$. Then, we aim to show the matrix identity:
$\begin{bmatrix}G_{f_{Q}}(t)&G_{f_{P}}(t)\\ \dot{G}_{f_{Q}}(t)&\dot{G}_{f_{P}}(t)\end{bmatrix}=\begin{bmatrix}G_{f_{Q}}(t-t_{j})G_{f_{Q}}(t_{j})+G_{f_{P}}(t-t_{j})\dot{G}_{f_{Q}}(t_{j})&G_{f_{Q}}(t-t_{j})G_{f_{P}}(t_{j})+G_{f_{P}}(t-t_{j})\dot{G}_{f_{P}}(t_{j})\\ \dot{G}_{f_{Q}}(t-t_{j})G_{f_{Q}}(t_{j})+\dot{G}_{f_{P}}(t-t_{j})\dot{G}_{f_{Q}}(t_{j})&\dot{G}_{f_{Q}}(t-t_{j})G_{f_{P}}(t_{j})+\dot{G}_{f_{P}}(t-t_{j})\dot{G}_{f_{P}}(t_{j})\end{bmatrix}.$ (67)
The right-hand side can be simplified using trigonometric identities, such as
$\cos(\Omega t)=\cos[\Omega(t-t_{j})]\cos(\Omega t_{j})-\sin[\Omega(t-t_{j})]\sin(\Omega t_{j}),$ (68)
$\sin[\Omega(t-t_{l+1})]=\cos[\Omega(t-t_{j})]\sin[\Omega(t_{j}-t_{l+1})]+\sin[\Omega(t-t_{j})]\cos[\Omega(t_{j}-t_{l+1})].$ (69)
By applying these identities, one can verify the equality of the $(1,1)$ and $(1,2)$ entries of the matrix in Eq. (67). The equalities of the $(2,1)$ and $(2,2)$ components then follow by taking the $t$-derivatives of the corresponding $(1,1)$ and $(1,2)$ entries.
Thus, the equality in Eq. (67) holds to first order in $\gamma/\Omega$.
Similarly, the noise matrix satisfies the composition law $\mathbb{N}_{t}=\mathbb{T}_{t-s}\mathbb{N}_{s}\mathbb{T}_{t-s}^{\intercal}+\mathbb{N}_{t-s}$, which can be shown by employing the same trigonometric identities as in the proof above.

### v3

@@ D.1 Transformation of covariance matrices
As described in Sec. IV.2, Gaussian states are fully determined by the first moment $\bar{\bm{R}}\equiv\braket{\bm{R}}$ and the covariance matrix $\mathbb{V}$. Any tensor product of Gaussians is Gaussian, and Gaussian operations are those CP maps preserving the Gaussianity. In particular, the full Hamiltonian in the Caldeira-Leggett model generates a Gaussian unitary. If the initial joint state is a product of Gaussian states, $\rho_{\text{tot}}(0)=\rho_{\text{s}}(0)\otimes\rho_{\text{E}}(0)$, then the final reduced state of the system, $\rho_{\text{s}}(t)=\operatorname{Tr}_{\text{E}}[U_{t}\rho_{\text{tot}}(0)U_{t}^{\dagger}]$, after the Gaussian unitary time-evolution $U_{t}\equiv U(t,0)$ remains Gaussian.
@@
D.2 Transformation of covariance matrices under the Markovian dynamics
Next, we consider how the above Markovianity condition (86) applies to the formulation using covariance matrices in Gaussian states. In particular, we show that if the time-evolution is described by a Markovian dynamical map, then the matrices $\mathbb{T}$ and $\mathbb{N}$ satisfy
@@ D.2 Transformation of covariance matrices under the Markovian dynamics
where $\mathbb{T}_{t-s}\equiv\mathbb{T}(t,s)$ and $\mathbb{N}_{t-s}\equiv\mathbb{N}(t,s)$ are the matrices associated with a CPTP map $\mathcal{E}_{t,s}$.
@@
where $\mathcal{E}_{t,0}$ is a CPTP map and we used the relation (63). Suppose the CPTP map $\mathcal{E}_{t,0}$ is a Markovian dynamical map. Then, we have the following relations:
@@
Appendix E Proof of Markovianity in Caldeira-Leggett model
In Appendix B, we showed that the Markovianity is realized when the memory kernel $\Sigma(t,t^{\prime})$ is restricted to nearest-neighbor propagation. In this section, we present an analogous analysis in the Caldeira–Leggett model.
@@ Appendix E Proof of Markovianity in Caldeira-Leggett model
$\mathbb{T}_{t}$ $=\begin{bmatrix}G_{f_{Q}}(t)&G_{f_{P}}(t)\\ \dot{G}_{f_{Q}}(t)&\dot{G}_{f_{P}}(t)\end{bmatrix}\,$, (112a)
$\mathbb{N}_{t}$ $=\begin{bmatrix}N_{QQ}(t)&N_{QP}(t)\\ N_{PQ}(t)&N_{PP}(t)\end{bmatrix}\,$, (112b)
where the functions $G_{f_{Q}}(t)$ and $G_{f_{P}}(t)$ are defined in (56), and the matrix components $N_{QQ}(t)$, $N_{QP}(t)$, $N_{PQ}(t)$, and $N_{PP}(t)$ are specified in (68).
Next, we consider transformation matrices (112) composed only of the neighboring propagation $\Sigma(t_{l+1},t_{l})$. By utilizing the trigonometric identities,
$\sin[\Omega(t-t_{j})]$ $=\sin(\Omega t)\cos(\Omega t_{j})-\cos(\Omega t)\sin(\Omega t_{j})\,$, (113a)
$\cos[\Omega(t-t_{j})]$ $=\cos(\Omega t)\cos(\Omega t_{j})+\sin(\Omega t)\sin(\Omega t_{j})\,$, (113b)
we find
$\mathbb{T}^{(\mathrm{M})}_{t}$ $=\bm{F}_{t}+\bm{F}_{t}\sum_{i_{1}=1}^{N-1}\bm{P}_{i_{1}}+\bm{F}_{t}\sum_{i_{1}=1}^{N-1}\sum_{i_{2}=1}^{i_{1}-1}\bm{P}_{i_{1}}\bm{P}_{i_{2}}+\cdots$
$=\bm{F}_{t}\,\mathsf{T}\exp\left(\sum_{j=1}^{N-1}\bm{P}_{j}\right)\,$, (114a)
$\mathbb{N}^{(\mathrm{M})}_{t}$ $=\sum_{k,k^{\prime}=1}^{N}\bm{F}_{t}\,\mathsf{T}\exp\left(\sum_{j=1}^{N-1}\bm{P}_{j}\right)\bm{g}(t_{k})\nu_{k,k^{\prime}}$
$\qquad\times\left[\bm{F}_{t}\,\mathsf{T}\exp\left(\sum_{j=1}^{N-1}\bm{P}_{j}\right)\bm{g}(t_{k^{\prime}})\right]^{\intercal}\,.$ (114b)
Here, $\bm{F}_{t,t^{\prime}}\coloneqq\bm{F}(t,t^{\prime})$ is a $2\times 2$ matrix
$\bm{F}_{t,t^{\prime}}$ $=\begin{bmatrix}\cos[\Omega(t-t^{\prime})]&\sin[\Omega(t-t^{\prime})]/\Omega\\ -\Omega\sin[\Omega(t-t^{\prime})]&\cos[\Omega(t-t^{\prime})]\end{bmatrix}\,$,
representing the unitary evolution of the system, $\bm{g}(t)=[-\sin(\Omega t)/\Omega,\cos(\Omega t)]^{\intercal}$ is a vector constructed from free Green’s function for a quantum harmonic oscillator, and
$\bm{P}_{j}$ $=\begin{bmatrix}-\sin(\Omega t_{j+1})X_{j}\cos(\Omega t_{j})/\Omega&-\sin(\Omega t_{j+1})X_{j}\sin(\Omega t_{j})/\Omega^{2}\\ \cos(\Omega t_{j+1})X_{j}\cos(\Omega t_{j})&\cos(\Omega t_{j+1})X_{j}\sin(\Omega t_{j})/\Omega\end{bmatrix}\,$, (115)
with $X_{j}=-\Sigma(t_{j+1},t_{j})T^{2}/N^{2}$.
For any integer $a\in(1,N)$ and $t=t_{N}$, the Markovianity condition (110) can be shown as
$\mathbb{T}_{t,t_{a}}\mathbb{T}_{t_{a}}$ $=\bm{F}_{t,t_{a}}\bm{F}_{t_{a}}\mathsf{T}\exp\left(\sum_{j=a}^{N-1}\bm{P}_{j}\right)\,\mathsf{T}\exp\left(\sum_{j=1}^{a-1}\bm{P}_{j}\right)\,$
$=\bm{F}_{t}\,\mathsf{T}\exp\left(\sum_{j=1}^{N-1}\bm{P}_{j}\right)=\mathbb{T}_{t}\,$, (116)
where we used the trigonometric identities (113) in the second equality. From a similar calculation, one can show that the noise matrix $\mathbb{N}_{t}^{(\mathrm{M})}$ satisfies the composition law (110).